# FunDiff: Diffusion Models over Function Spaces for Physics-Informed Generative Modeling

### v1

@@ 3.3 Methods
To realize the framework of function generative models outlined in Section 3.1, we propose FunDiff (Fig. 1), which consists of a novel FAE and a latent diffusion model to generate continuous functions. The FAE learns the mapping from discrete function representations to continuous ones, while also enabling the enforcement of prior physical constraints. The latent diffusion model is then trained directly on the latent manifold produced by the pretrained FAE encoder. Our generation process follows a simple two-step procedure: we first sample latent codes by integrating the reverse-time ODE from Gaussian noise to the target data distribution, and then transform these codes into continuous functions via the FAE decoder.
3.3.1 Network architecture
@@
The encoder $\mathcal{E}$ employs a ViT backbone to extract features and a Perceiver module to handle functions of different discretizations. Input functions $\mathbf{f}\in\mathbb{R}^{H\times W\times C}$ are first processed by ViT-style patchification, resulting in flattened patches $\mathbf{f}_{p}\in\mathbb{R}^{L\times D}$, where $L=\frac{H}{P}\times\frac{W}{P}$ represents the length of the sequence, $P$ is the size of the patch, and $D$ denotes the embedding dimension. To accommodate varying resolutions, trainable positional embeddings (PE) are dynamically interpolated to match the input sequence length $L$, yielding position-aware patch embeddings $\mathbf{f}_{pe}=\mathbf{f}_{p}+$ Interp$(\mathrm{PE},L)$.
@@
The core of our decoder uses the embedded coordinates $\mathbf{x}_{0}\in{\mathbb{R}}^{D}$ as queries, with the encoder output $\mathbf{z}_{L}$ serving as both keys and values in a series of $K$ cross-attention Transformer blocks:
@@
Physical prior integration.
$\mathcal{L}_{\text{PDE}}=|\mathcal{R}(\hat{f})|_{2}^{2},$
where $\mathcal{R}$ represents the PDE residual and $\hat{f}$ is the reconstructed function. The continuous representation of FAE enables automatic differentiation to compute all required derivatives, allowing constraint enforcement across the entire domain.
Furthermore, some physical properties can be incorporated through architectural modifications. For periodic boundary conditions, one approach is to modify the coordinate embedding with Fourier bases of specific frequencies. When handling divergence-free fields, particularly important in fluid dynamics applications, the FAE can be configured to learn a stream function $\psi$, from which velocity field components are derived as $u=\frac{\partial\psi}{\partial y},\quad v=-\frac{\partial\psi}{\partial x}$, naturally ensuring divergence-free flow. Symmetry constraints can be incorporated when needed through modifications to the decoder’s final layer.
Here, we emphasize that our approach fundamentally differs from existing physics-informed diffusion models [37, 38, 39, 40, 32], which impose physical constraints either by modifying the diffusion model’s training objective or by guiding the posterior sampling process. These methods operate on discretized outputs, where physical laws—typically enforced through PDE residuals—are subject to discretization errors. This not only leads to inaccurate constraint enforcement but also introduces additional optimization challenges. Moreover, in inverse problems where key physical variables are unobserved, such residuals may be undefined at inference time, further limiting the applicability of these approaches. In contrast, our method enforces physical constraints within the FAE, which is fully decoupled from the training and inference processes of the latent diffusion model. Since the FAE provides a continuous neural representation, we can compute PDE residuals via automatic differentiation when needed. This design enables both exact and soft enforcement of constraints in continuous space, mitigating discretization-induced errors and enabling more stable and physically consistent generation. Details on imposing these constraints are provided in Section 4 through comprehensive numerical experiments, demonstrating that our framework can generate diverse function samples while naturally preserving the underlying physical principles.
@@
The DiT backbone comprises a sequence of Transformer blocks that utilize AdaLN-Zero modulation [58], an enhanced variant of Adaptive Layer Normalization designed for stable time-dependent conditioning. For each block, the modulation parameters are computed as
@@ 3.3.2 Training
[…]
@@ 3.3.3 Inference
starting from a random Gaussian noise $\mathbf{z}_{1}\sim\mathcal{N}(0,\mathbf{I})$ and integrating backward to obtain $\mathbf{z}_{0}$. This process is deterministic and typically requires 20–50 function evaluations for high-quality samples. Second, we decode the sampled latent codes through the FAE decoder to obtain the final continuous function. Importantly, since our decoder operates on arbitrary query points, we can evaluate the generated function at any desired resolution or location in the domain.
4 Results
In this section, we demonstrate the effectiveness of our proposed FunDiff framework through a series of experiments on diverse physical systems. Specifically, we present four case studies with increasing complexity: (1) the Burgers’ equation, showcasing the incorporation of PDE constraints; (2) Kolmogorov flow, highlighting divergence-free field generation; (3) linear elasticity, illustrating symmetry preservation; and (4) turbulent mass transfer, demonstrating conditional generation capabilities.
@@ 4 Results
For the FAE training, we employ the AdamW optimizer [59] with a weight decay of $10^{-5}$. The learning rate follows a two-phase schedule: an initial linear warm-up from zero to $10^{-3}$ over 2,000 steps, followed by exponential decay with a factor of 0.9 every 2,000 steps. Each training batch contains 16 samples, with 4,096 query coordinates and their corresponding outputs randomly sampled from the grid. To enhance model’s resolution invariance, we randomly downsample inputs by different factors (e.g., 1, 2, 4, 8) during training while optimizing the model to reconstruct the original high-resolution input function by minimizing the MSE loss.

### v2

@@ Results
Learning 1D damped sinusoidal functions
Figure 2: Damped sinusoidal functions. Analysis of generative quality across models trained with different numbers of samples ($n=16,32,64,128,256$). (Top) Comparison between the worst-case generated samples (blue dots) and their reconstructed signals (red dashed lines) obtained through parameter estimation. (Bottom) The first four panels show the estimated parameter distributions for amplitude (AA), damping coefficient ($\gamma$), angular frequency ($\omega$), and vertical shift ($b$). Red dashed lines indicate the true parameter bounds from the underlying uniform distributions. The rightmost panel displays the MSE of the reconstructed signal. The results demonstrate that models trained with more samples generate functions that better align with the true distributions and achieve lower reconstruction error.
$f(x)=A\exp(-\gamma x)\sin(\omega x)+b,\quad x\in[0,1],$ (2)
We employ an FAE with a 6-layer encoder and a 2-layer decoder, each with a hidden dimension of 256 and 8 attention heads. The FAEs are trained on datasets of varying sizes, with function samples discretized at 128 uniform grid points. We train the FAE by minimizing the loss using the Adam optimizer for $5\times 10^{4}$ iterations. The learning rate follows a warmup and exponential decay schedule: it linearly increases from 0 to $10^{-3}$ over the first 1,000 steps, then exponentially decays every 1,000 steps with a decay rate of 0.9. Following the FAE training, we train the latent diffusion model by minimizing (9) using an identical learning rate schedule.
For evaluation, we generate 8,192 sample functions using our inference procedure with 100 denoising steps. We assess the quality of these generated samples through parameter estimation, leveraging the analytical form of damped sinusoidal functions. Specifically, for each generated sample, we perform a least-squares fitting using (2) to obtain the inferred parameters and reconstruct the corresponding signal. This approach enables a rigorous quantitative assessment of our model’s ability to capture the underlying function space through two metrics: the comparison between reconstructed and generated signals, and the analysis of inferred parameter distributions against the target uniform distributions.
We provide a comprehensive evaluation of our model’s generative performance across varying training dataset sizes (Fig. 2). The top row illustrates the worst-case generated samples alongside their reconstructions via parameter estimation. As the number of training samples increases, the quality of generation improves significantly. The bottom row analyzes the parameter distributions, revealing the evolution of our model’s capacity to capture the underlying function space. With limited training data (16 or 32 samples), the estimated parameter distributions exhibit narrower ranges or deviate significantly from the target uniform distributions, indicating insufficient coverage of the parameter space. As the size of training dataset increases, the estimated parameter distributions progressively converge to the target uniform distributions, with the model trained on 256 samples achieving notably better alignment with the true distributions. This improvement is quantitatively validated by the decreasing mean squared error (MSE) between the generated samples and their reconstructions. These empirical results strongly support our theoretical analysis.
Kolmogorov flow
In this example, we use the proposed FunDiff to generate divergence-free velocity fields, a fundamental requirement for ensuring the physical consistency of incompressible flow simulations in computational fluid dynamics. We consider two scenarios: unconditional flow generation and flow reconstruction from partial observations. To evaluate our approach, we use the two-dimensional Kolmogorov flow, a well-established benchmark for analyzing transition to turbulence and the stability of incompressible fluid systems. The governing equations are formulated in the vorticity-velocity representation as follows:
$\frac{\partial\omega(x,t)}{\partial t}+\mathbf{u}(x,t)\cdot\nabla\omega(x,t)$ $=\frac{1}{\operatorname{Re}}\nabla^{2}\omega(x,t)+f(x),$
$\nabla\cdot\mathbf{u}(x,t)$ $=0,$
$\omega(x,0)$ $=\omega_{0}(x),$
where $\omega$ is the vorticity, $\mathbf{u}=(u,v)$ is the velocity field, and $\operatorname{Re}$ is the Reynolds number, set to 1000. The forcing term $f(x)$ drives the flow and $x=[x_{1},x_{2}]$. The equations are defined with periodic boundary conditions in the spatial domain $x\in(0,2\pi)^{2}$ and time interval $t\in(0,T]$. The forcing term is defined as
$f(x)=-4\cos\left(4x_{2}\right)-0.1\omega(x,t),$
where the additional drag force term $0.1\omega(x,t)$ prevents energy accumulation on large scales.
$\left[x,y,\mathcal{E}_{\theta}(f)\right]\xrightarrow{\mathcal{D}_{\theta}}\psi(\gamma(x,y)),$ (3)
$\gamma(x,y)=[\cos(x),\sin(x),\cos(y),\sin(y)].$ (4)
This embedding guarantees that the output functions have a periodicity of $2\pi$ within $0,2\pi]^{2}$. An incompressible velocity field is then computed as $(u,v)=(\partial\psi/\partial y,-\partial\psi/\partial x)$.
Unconditional flow generation
Following the training procedure in the Method section, we present our results in Fig. 3a, which displays representative generated samples along with their corresponding divergence fields, computed via automatic differentiation. The predicted divergence values are negligible, confirming the effectiveness of our approach. We also compare the spectral energy distributions between the training dataset and 1,000 generated samples (Fig. 3b). The spectra show excellent agreement up to wavenumbers of $\mathcal{O}(10^{2})$. However, the model leads to some artificial high-frequency components at larger wavenumbers.
This unconstrained approach results in significantly higher divergence values (Fig. S1). Additionally, the spectral energy shows some discrepancies compared to the model constrained by the exact divergence-free condition.
Flow reconstruction from partial observations
Furthermore, we benchmark our approach against several state-of-the-art neural operators and PDE surrogate models, including FNO [13], ViT [8], and UNet [60], as well as more recent architectures such as CViT [61], DPOT [62], CNextU-Net [63], and AViT [64]. To further evaluate performance under inverse problem settings, we also include three diffusion-based baselines: Diffusion Posterior Sampling (DPS) [65], DiffusionPDE [42], and Spatially-Aware Diffusion [51]. Detailed implementation specifications for these baseline models are provided in Appendix S4.1. Fig. 3c presents reconstructed flow fields from various models, where the input consists of 64 × 64 resolution velocity fields corrupted by 20% Gaussian noise. It can be observed that neural operator baselines tend to produce highly discontinuous or spatially inconsistent patterns, whereas diffusion-based models recover coherent large-scale structures but often lose fine-scale features. The results demonstrate that our model uniquely achieves both smooth and accurate reconstructions. Additional visualizations of our model’s predictions, with and without the divergence-free condition imposed in FAE decoder, are shown in Figs. S2 and S3, respectively. The quantitative comparison using relative $L^{2}$ errors across different downsampling factors is presented in Fig. 3e. These results confirm that our FunDiff model maintains consistent performance across all tested resolutions while exhibiting strong resilience to noise.
Burgers’ equation
As our next example, we demonstrate the effectiveness and flexibility of our framework in generating physically consistent samples through the incorporation of physical laws as soft constraints in the training of FAE. To this end, we consider the Burgers’ equation, which serves as a fundamental prototype for conservation laws that can develop discontinuities. The governing equation is given by
$\frac{du}{dt}+u\frac{du}{dx}-\nu\frac{d^{2}u}{dx^{2}}=0,\quad(x,t)\in(0,1)\times(0,1],$ (5)
$u(x,0)=a(x),\quad x\in(0,1)$ (6)
with periodic boundary conditions, where $\nu=0.001$ denotes the viscosity coefficient, and the initial condition $a(x)$ is sampled from a Gaussian random field (GRF) satisfying periodic boundary conditions.
Let $u_{\theta}$ denote the FAE output. We define the PDE residual as
The FAE is trained by minimizing a composite loss
$\mathcal{L}(\mathbf{\theta}):=\underbrace{w_{data}\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\left|{u}_{\mathbf{\theta}}^{i}(x_{j},t_{j})-{u}^{i}(x_{j},t_{j})\right|^{2}}_{\mathcal{L}_{data}(\mathbf{\theta})}+\underbrace{w_{physics}\frac{1}{NM}\sum_{i=1}^{N}\sum_{j=1}^{M}\left|\mathcal{R}[{u}_{{\theta}}^{i}](x_{j},t_{j})\right|^{2}}_{\mathcal{L}_{physics}(\mathbf{\theta})}.\$
Similar to the Kolmogorov flow example, our objective here is to train a FunDiff model to reconstruct the solution field from sparse observations. We evaluate our model against established baseline methods. We show three examples in Fig. 4a. Neural operator baselines generally produce blurry and discontinuous reconstructions, whereas diffusion-based approaches yield smoother results but often exhibit small oscillations near sharp gradients. In contrast, our FunDiff generates sharp and smooth reconstructions, which are in excellent agreement with the reference numerical solutions. Fig. 4b presents a quantitative comparison of model performance across inputs with varying downsampling factors and 20% Gaussian noise corruption. Notably, FunDiff outperforms all baselines. Additional examples of our model’s predictions are shown in Fig. S4. Their physical consistency is confirmed by low PDE residuals, remaining on the order of $\mathcal{O}(10^{-2})$ across the entire domain. We also systematically examine the effect of noise and data sparsity (Fig. 4c). As expected, reconstruction error increases with higher noise levels, but FunDiff demonstrates strong robustness to data sparsity, maintaining stable performance even under aggressive downsampling (factors up to 5).
To rigorously evaluate the impact of physical constraints, we conduct an ablation study by training the function encoder without the physics loss term $\mathcal{L}_{physics}$. The absence of PDE-based regularization leads to significantly higher predictive errors and PDE residuals in the generated solutions (Fig. 4d). This degradation in physical consistency is further visualized through the generated samples in Fig. S5, underscoring the crucial role of physics-informed training in our framework.
Linear elasticity
Symmetries play an important role in physics, providing a framework for understanding conservation laws and invariances that govern natural phenomena. This example aims to demonstrate how to embed prior symmetries into FunDiff models. To this end, we consider linear elasticity in computational micromechanics, which involves solving the Lippmann-Schwinger equation [66] in a 2D periodic representative volume element (RVE) microstructural domain $\Omega\subset\mathbb{R}^{2}$ with spatial coordinates $\mathbf{x}=(x,y)$ under plane strain conditions. The problem is formulated as
[…]
These relations represent balance and constitutive equations, respectively, which are written in a component-wise manner ($1\leq i,j,k,l\leq 2$) using the Voigt notation. $\mathbf{x}$ and $\mathbf{x}^{\prime}$ denote a pair of target and source points inside the RVE domain $\Omega$. $\sigma_{ij},\varepsilon_{ij}$ and $C_{ijkl}$ denote stress, strain, and fourth-order elastic tensor, respectively. $C_{ijkl}(\mathbf{x})$ is expressed in terms of local Young’s modulus $E(\mathbf{x})$ and Poisson ratio $\nu(\mathbf{x})$ as
where $\delta_{ij}$ denotes the Kronecker delta function. $G_{ijkl}^{(0)}$ denotes Green’s operator corresponding to the homogeneous reference material with elastic tensor $C_{ijkl}^{0}$ [67]. $\bar{\varepsilon}_{ij}$ denotes far field strain used to apply the strain-controlled boundary condition, equal to the average strain over the RVE domain $\Omega$. In this example, we model the point-wise strain concentration tensor $A_{ijkl}(\mathbf{x})\in\mathbb{R}^{3\times 3}$, which linearly maps $\bar{\varepsilon}_{ij}$ to $\varepsilon_{ij}(\mathbf{x})$ independent of loading conditions:
$A_{ijkl}(\mathbf{x})$ can be computed by triply solving (7) subject to three orthogonal unit far field strain $\overline{\varepsilon}_{ij}$ as
$\overline{\varepsilon}_{ij}=\begin{Bmatrix}[1,0,0]^{T},[0,1,0]^{T},[0,0,1]^{T}\end{Bmatrix}.$ (8)
Due to the intrinsic symmetry of the RVE configuration, the resulting strain concentration tensor field $A_{ijkl}(\mathbf{x})$ exhibits a symmetric pattern with respect to both $x=\frac{1}{2}$ and $y=\frac{1}{2}$. We aim to train a FunDiff model to reconstruct the strain concentration tensor field from sparse observations while respecting the inherent symmetry priors. To this end. we modify the forward pass of the decoder to enforce these constraints. Let $\mathbf{z}$ denote the FAE encoder output latents. The symmetry-preserved decoder $\hat{\mathcal{D}}_{\theta}$ is then constructed as
Fig. 5a presents reconstructed samples generated by our FunDiff model from a coarse strain concentration field corrupted by 10% Gaussian noise. The results show that our model preserves underlying symmetries and accurately resolves sharp material interfaces. To assess robustness against different types of noise, we report test errors in Fig. 5b. The consistently low errors across all noise configurations highlight the framework’s reliability in practical applications.
To further demonstrate the critical role of the continuous decoder in our framework, we implement a baseline using a discrete autoencoder with a Vision Transformer encoder and an MLP decoder projecting to the target pixel space. We train a diffusion transformer under identical conditions, with representative generated samples shown in Fig. S6. Unlike our approach, the baseline fails to preserve symmetry or enforce physical continuity in strain fields, instead producing piecewise discontinuous predictions.
Turbulence mass transfer
In the last example, we examine the model capacity of inferring non-intrusive quantities such as pressure in turbulent problems. Specifically, we consider turbulent mass transfer in a rectangular channel. This phenomenon plays a critical role in the chemical engineering process, directly influencing reaction kinetics, conversion efficiency, and product purity. Our study focuses on a computational domain $\Omega=\{(x,y)\mid 0.006<x<0.018,0<y<0.01\}$. Within this domain, we are interested in training a FunDiff model to infer the obstacle topology and pressure field given the sparse and noisy velocity measurements.
$\frac{\partial(\rho u_{i})}{\partial x_{i}}=0,\quad$
$u_{i}\frac{\partial(\rho u_{j})}{\partial x_{i}}=-\frac{\partial P}{\partial x_{j}}+\frac{\partial}{\partial x_{i}}\left[\left(\mu_{m}+\mu_{t}\right)\frac{\partial u_{j}}{\partial x_{i}}\right],\quad$
$u_{i}\frac{\partial(\rho C)}{\partial x_{i}}=\frac{\partial}{\partial x_{i}}\left[\rho\left(D_{m}+D_{t}\right)\frac{\partial C}{\partial x_{i}}\right],\quad$
where $p,u_{i},x_{i}$ are the pressure, velocity and location respectively; the subscript $i$ represents the coordinate direction which can take $x$ and $y$ for the two-dimensional system. Here, $C$ is the time-averaged concentration; $\rho,\mu_{m}$ and $D_{m}$ are density, laminar viscosity, and laminar mass diffusivity, which are physical property constants and fixed to $1000\mathrm{\penalty 10000\ kg}/\mathrm{m}^{3},1.003\times 10^{-6}\mathrm{\penalty 10000\ m}^{2}/\mathrm{s},1.003\times 10^{-8}\mathrm{\penalty 10000\ m}^{2}/\mathrm{s}$, respectively. $\mu_{t}$ and $D_{t}$ are the turbulent viscosity and turbulent mass diffusivity, respectively.
Representative generated samples (Fig. 6a) validate the model’s ability to infer pressure fields and reconstruct obstacle geometries despite heavily corrupted input velocity measurements. The reconstructions align closely with numerical estimations. Fig. 6b presents the test error across various downsampling factors, noise levels, and noise types. The results indicate that our model remains robust even under severe noise corruption and low-resolution inputs. Our results demonstrate the model’s effectiveness in preserving critical flow features and geometric details, and highlight the framework’s potential for real-world applications where high-fidelity reconstructions are required from sparse or noisy measurements.
Discussion
In this work, we propose FunDiff, a flexible and efficient framework for generative modeling in function spaces, addressing fundamental challenges in generating continuous function data while preserving physical constraints. Our primary contributions are twofold: (1) a rigorous theoretical foundation establishing conditions for function distribution approximation via autoencoders, and (2) a novel constrained function autoencoder (FAE) that inherently enforces physical principles. Through extensive experimentation across diverse physical systems, we have demonstrated that our approach successfully generates physically valid samples while maintaining high fidelity to the target PDE solution distributions.
While our current work lays a strong foundation for physics-informed generative modeling, several promising directions for future research emerge. One particularly interesting direction is to develop not purely data-driven, but physics-informed diffusion-based PDE solvers. This could be achieved by incorporating PDE residuals directly into the score function of diffusion models and by using additional initial and boundary conditions as guidance during sampling, thereby steering the generated samples toward the corresponding PDE solutions. Another natural extension would be to adapt our framework to handle more complex geometries, thereby enabling broader applicability in real-world scientific and engineering problems. Handling complex domains necessitates incorporating geometric inductive biases, such as mesh-based representations, implicit neural surfaces, or graph-based function approximations. Moreover, extending our approach to multiphysics applications represents a critical and exciting frontier. Many engineering and scientific challenges involve the interplay of multiple physical processes, such as fluid-structure interactions, thermo-mechanical coupling, and electromagnetic-thermal phenomena, which require models that can simultaneously capture and reconcile diverse physical constraints. These advancements will significantly expand the scope of generative modeling for physical systems and ultimately contribute to more accurate and efficient simulation-based design across disciplines.
Methods
To realize our theoretical framework for generative modeling in function spaces, we develop FunDiff, a practical implementation that addresses the key challenges of handling infinite-dimensional function distributions while preserving physical constraints. Our approach must accommodate discretized functions at varying resolutions, enable continuous evaluation at arbitrary coordinates, and flexibly incorporate domain-specific physical priors without compromising the generative modeling process. The architecture is designed to decouple physical constraint enforcement from diffusion training, ensuring stable optimization while maintaining mathematical rigor in constraint satisfaction. We detail the FunDiff network architecture, training procedures, and constraint integration mechanisms that enable robust function generation across diverse scientific domains.
Network architecture
@@ Methods
The encoder $\mathcal{E}$ employs a ViT backbone to extract features and a Perceiver module to handle functions of different discretizations. Input functions $\mathbf{f}\in\mathbb{R}^{H\times W\times C}$ are first processed by ViT-style patchification, resulting in flattened patches $\mathbf{f}_{p}\in\mathbb{R}^{L\times D}$, where $L=\frac{H}{P}\times\frac{W}{P}$ represents the length of sequence, $P$ is the size of the patch, and $D$ denotes the embedding dimension. To accommodate varying resolutions, trainable positional embeddings (PE) are dynamically interpolated to match the input sequence length $L$, yielding position-aware patch embeddings $\mathbf{f}_{pe}=\mathbf{f}_{p}+$ Interp(PE,L)($\mathrm${PE},L).
@@
The core of our decoder uses embedded coordinates $\mathbf{x}_{0}\in{\mathbb{R}}^{D}$ as queries, with the encoder output $\mathbf{z}_{L}$ serving as both keys and values in a series of $K$ cross-attention Transformer blocks:
@@
The DiT backbone consists of a sequence of Transformer blocks that utilize AdaLN-Zero modulation [76], an enhanced variant of Adaptive Layer Normalization designed for stable time-dependent conditioning. For each block, the modulation parameters are computed as
@@
$\mathcal{L}_{\text{DiT}}$ $=\int_{0}^{1}\mathbb{E}_{\mathbf{f}\sim p_{\text{data}}}\left[\left|\left(\mathbf{z}_{1}-\mathbf{z}_{0}\right)-\mathbf{g}(\mathbf{z}_{\tau},\tau)\right|^{2}\right]d\tau,\$$\text{with}\quad\mathbf{z}_{\tau}=(1-\tau)\mathbf{z}_{1}+\tau\mathbf{z}_{0},\quad\mathbf{z}_{1}=\mathcal{E}(\mathbf{f}),$ (9)
@@
starting from a sample of standard Gaussian noise $\mathbf{z}_{1}\sim\mathcal{N}(0,\mathbf{I})$ and integrating backward to obtain $\mathbf{z}_{0}$. This process is deterministic and typically requires 20–50 function evaluations for high-quality samples. Second, we decode the sampled latent codes through the FAE decoder to obtain the final continuous function. Importantly, since our decoder operates on arbitrary query points, we can evaluate the generated function at any desired resolution or location in the domain.
@@
For the FAE training, we employ the AdamW optimizer [77] with a weight decay of $10^{-5}$. The learning rate follows a two-phase schedule: an initial linear warm-up from zero to $10^{-3}$ over 2,000 steps, followed by exponential decay with a factor of 0.9 every 2,000 steps. Each training batch contains 16 samples, with 4,096 query coordinates and their corresponding outputs randomly sampled from the grid. To enhance the model’s resolution invariance, we randomly downsample inputs by different factors (e.g., 1, 2, 4, 8) during training while optimizing the model to reconstruct the original high-resolution input function by minimizing the MSE loss.